OpenStudy (anonymous):

what is the simplified form of 6√x*6√x*6√x*6√x

OpenStudy (anonymous):

is that \[6\sqrt{x}\times6\sqrt{x}\times6\sqrt{x}\times6\sqrt{x}\]